OpenStudy (lgbasallote):

how many three letter words can be formed from LGBASALLOTE? note: words not necessary to be in the dictionary

OpenStudy (lgbasallote):

is it \[\large \frac{11!}{8!}\]

OpenStudy (callisto):

11x10x9 , maybe

OpenStudy (accessdenied):

You'd divide off each individual factorial for the number of repeated letters Like there are 3 L's so divide off 3! There are 2 A's so divide off 2! as well etc although i guess it depends on how you define a word, just as a group of letters or has to have a vowel or something. D:

OpenStudy (anonymous):

it is 11!/8!. I just did a very similar problem. The equation to use is (n!)/(n-r)!

OpenStudy (lgbasallote):

but that's the number of how many times LGBASALLOTE can be arranged @AccessDenied

OpenStudy (lgbasallote):

thanks :D

OpenStudy (accessdenied):

Hmm... where does the 8! come from?

OpenStudy (lgbasallote):

11 - 3

OpenStudy (accessdenied):

oh wait formula ic

OpenStudy (callisto):

Hmm... 11! / 8! = 11x10x9 Since I'm not good at using ''!'' I usually just think it in a VERY traditional way... there are 11 letters to choose for the 1st letter of the word, 10 letters to choose for the 2nd letter of the word, 9 letters to choose for the 3rd letter of the word So, multiply them all... And get the answer *I wish someone can teach me these stuffs :( *

OpenStudy (accessdenied):

OH okay. I see where this is coming from. D: My mind just read problem and said "how many ways can you arrange it uniquely" automatically lol.

OpenStudy (lgbasallote):

i hear you girl..im also trying to self-learn these stuffs @_@

OpenStudy (accessdenied):

Hmm.. but then would you have to consider non-unique cases, like choosing ASA and then the other A first for ASA?

OpenStudy (lgbasallote):

that...is a good point

OpenStudy (callisto):

hmm? That has been taken into account already?! Since I haven't specified which one to choose ...

OpenStudy (lgbasallote):

maybe it should be \[\large \frac{11!}{8!3!2!}\]

OpenStudy (accessdenied):

like, in your case you consider all letters as unique; like, you choose A or the other A it's considered as different cases But if you consider ASA and ASA as the same word, you'd have to eliminate the repeats. D:

OpenStudy (anonymous):

This gives a lot of really useful information - http://www.mathsisfun.com/combinatorics/combinations-permutations.html

OpenStudy (callisto):

My English fails me.... :S

OpenStudy (lgbasallote):

so \[\frac{6!}{8!}\] wouldnt make sense >.<

OpenStudy (anonymous):

None of the answer posted is correct :)

OpenStudy (callisto):

No That does't make sense @lgbasallote

OpenStudy (callisto):

Wow~~ FFM comes to save us!!!

OpenStudy (accessdenied):

11! / 8! 3! 2! 11*9*8 ------ 6*2 11*3*2 ?

OpenStudy (anonymous):

Nopes, The word "L G B A S A L L O T E" has 3 L's and 2 A's the rest are all different.

OpenStudy (lgbasallote):

\[\large \frac{6!}{3!}\] seems few...

OpenStudy (anonymous):

There is no easy way to solve this one. I will do it with generating functions.

OpenStudy (lgbasallote):

oh..so this isnt simple :/ these things are hard to see if they are difficult or not..

OpenStudy (accessdenied):

wait i didn't even do my math correctly. blah. ;[ well, i guess its no use if it's not that anyways. lol

OpenStudy (anonymous):

If you guys want I can post the generating function approach.

OpenStudy (accessdenied):

I'm personally curious about it myself. :)

OpenStudy (callisto):

Go~~~

OpenStudy (anonymous):

The coefficient of \(x^3\) in: \[3! \times (1+x)^6 \times \left(1+x+\frac {x^2}{2!} \right )\times \left(1+x+\frac {x^2}{2!} +\frac{x^3}{3!}\right )\] If my algebra is correct the answer is \(379\)

OpenStudy (accessdenied):

I get 379 as well when I worked it out on paper. Very strange method to me, so I guess I have to study more into the subject later to understand better. D: Thanks, FoolForMath! :D

OpenStudy (anonymous):

Its' is actually very easy, if you are interested I will show you how to understand and derive it on your own.

OpenStudy (accessdenied):

Hmm, that'd be really interesting. :D Like, I can kind of see some patterns in the expression but not completely obvious how it all works out.

OpenStudy (anonymous):

First read about Generating function from here : http://users.softlab.ntua.gr/~nidal/discrete/notes/examples/gen2.pdf and here: http://courses.csail.mit.edu/6.042/fall05/ln11.pdf and then read this answer: http://math.stackexchange.com/questions/20238/

OpenStudy (anonymous):

Take it slow, and enjoy the ride :)